CAVITY RINGDOWN SPECTROSCOPY OF THE Ã^{2}A′ - X̃^{2}A′′ ELECTRONIC TRANSITION OF THE PHENYL PEROXY RADICAL

\underline{GABRIEL M.P. JUST}, ERIN N. SHARP, SERGEY J. ZALYUBOVSKY, AND TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Peroxy radicals are key intermediates in both combustion and atmospheric chemistry. We have previously studied small alkyl peroxy radicals (methyl, ethyl peroxy^{a}, 1- and 2-propyl peroxy^{b}) via their near IR electronic transition using cavity ringdown spectroscopy (CRDS). The reaction of phenyl radical (C_{6}H_{5}) with oxygen is postulated to play a particulary important role in the soot formation process inherent in hydrocarbon combustion. In this talk, we report CRDS of the near IR Ã^{2}A′-X̃^{2}A′′ electronic transition of phenyl peroxy radical. Ab initio calculations have been carried out to help assign T_{00} band origin, as well as other vibrational bands.

aM. B. Pushkarsky, S. J. Zalyubovsky, and T. A. Miller, J. Chem. Phys \underline{112}, 10695, (2000).

bS. Zalyubovsky, B. Glover, T. A. Miller, C. Hayes, J. K. Merle, and C. M. Hadad, J. Chem. Phys accepted